From Part A:

Mass of KxFe(C2O4)y · zH2O prepared : 4.200 g
Mass of FeCl3 : 1.60 g

From Part B:

% Potassium in compound : 22.00 %
% Iron (from ion exchange & titration vs. NaOH) : 9.10 %

From Part C:

% Oxlate : 36.96 %

- Calculate the % water of hydration :

Calculate the following for Fe3+:

g in 100 g sample mol in 100 g sample mol/mol Fe
(3 sig figs)
mol/mol Fe
(whole number)

Enter the simplest formula of the Iron Oxalate Complex Salt:

K Fe(C2O4) · H2O

Calculate the moles of FeCl3 used in preparation

Calculate the theoretical moles of KxFe(C2O4)y · zH2O:

Calculate the actual moles of KxFe(C2O4)y · zH2O synthesized

Calculate the percent yield.
Answer #1

%wt of H2O = 100 - (%wt of K + %wt of Fe + %wt of C2O4) = 31.94

Fe+3 in 100 grams of sample is 9.10 (given 9.1 percent)

number of moles of Fe+3 is (not clear in the question)

formulae of complex is K3[Fe(C2O4)3].10H2O

Number of moles of FeCl3 used in preparation of complex is 0.0099 (wt/M.wt = 1.6/162.5)

Theoritical moles of complex must be equal to number of moles of FeCl3 used = 0.0099

Actual moles of complex formed is 0.0068 (wt/M.wt = 4.2/617)

Percentage yield of complex is 68.6868